Circular No. 5902

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 S. Nakano, Sumoto, Japan, reports the discovery by Kazuyoshi Kana-

tsu, Matsue, Shimane, of a nova on a T-Max 400 exposure with a 55mm f/2.8 lens on Dec. 7, with confirmation by the discoverer on Dec.

11. K. Kawanishi, Akou, Hyogo, has determined the position to be α = 23^{h}39^{m}22^{s}.36, δ = +57^{o}14′23′′.7 (equinox 1950.0). Magnitude estimates: Dec. 5.44 UT, [10.0 (A. Tago, Tsuyama, Okayama; T-Max); 7.47, 7.5 (Tago); 7.47, 6.5 (Kanatsu); 11.42, 6.5 (Kanatsu); 11.63, 6.5 (S. Otomo, Kiyosato, Yamanashi; visual); 11.63, 6.4 (T. Urata, Shimizu, Shizuoka; unfiltered CCD); 11.65, 6.4 (K. Tomita, Tokyo; visual).

PERIODIC COMET MUELLER 5 (1993s)

Further precise positions have been reported as follows:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 10.40873 6^{h}51^{m}29.^{s}67 +21^{o}28′58′′.0  Scotti
10.41679 6 51 29.42 +21 28 59.3 17.4  _{”}
10.42585 6 51 29.13 +21 29 00.6 17.5  _{”}
10.44653 6 51 28.55 +21 29 08.3 18  Mueller

J. V. Scotti (Lunar and Planetary Laboratory). 0.91-m Spacewatch tele-

scope, Kitt Peak. m_{2} = 20.6. Coma diameter 8′′, 0′.43 tail in p.a. 312^{o}. J. Mueller, D. J. Griffith and J. D. Mendenhall (Palomar). 1.2-m Oschin

telescope. Comet less than 4 cm from edge of Sky Survey plate.

An orbit computation from the 15 observations Nov. 20-Dec. 10 confirms that the comet is of short period:

T = 1995 Feb. 8.153 TT  ω = 52.^{o}773
e = 0.33181  Ω = 100.572 }2000.0
q = 3.95249 AU  i = 17.104
a = 5.91527 AU n^{o} = 0.068508 P = 14.39 years
1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 9 6^{h}52.^{m}19 +21^{o}24.^{′}8 3.543 4.454 154.^{o}8 5.^{o}4 17.7
19 6 46.82 +21 55.6 3.471 4.433 166.3 3.0 17.7
29 6 40.72 +22 27.9 3.430 4.413 177.9 0.5 17.6
Jan. 8 6 34.41 +23 00.1 3.420 4.393 170.4 2.1 17.6
18 6 28.43 +23 31.2 3.442 4.373 158.8 4.7 17.6

1993 December 11 Brian G. Marsden